数学 I 年組名前

\begin{matrix}1 y = 2x^{2} + 4x + 3 のグラフをかけ． \\ x & y \\ 2 y = 2x^{2} − 8x + 9 のグラフをかけ． \\ x & y \end{matrix}  \begin{matrix}3 y = x^{2} − 8x + 13 のグラフをかけ． \\ y & x \\ 4 y = x^{2} − 8x + 13 のグラフをかけ． \\ y & x \end{matrix}

\begin{matrix}5 y = x^{2} + 6x + 13 のグラフをかけ． \\ x & y \\ 6 y = x^{2} + 2x − 4 のグラフをかけ． \\ y & x \end{matrix}  \begin{matrix}7 y = x^{2} + 2x − 1 のグラフをかけ． \\ x & y \\ 8 y = 2x^{2} + 8x + 9 のグラフをかけ． \\ x & y \end{matrix}